[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 01, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Oct 2 17:59:57 UTC 2016


HSU Shift Log for Sat Oct 01 2016 22:54:21 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS/AWIPS2 issues:

Other Techincal issues: Satellite data went down after 23Z- NESDIS software  
issue. We formally requested the GOES-W satellite to be put into backup  
mode (30 min full disk) at 0135Z since we have no plane until 1130Z. Five  
minutes later, the satellite data came back and appears to be working  
normally, and the backup plan was apparently discontinued by the SDM.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161002/4147bf58/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list